• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Aniol

    Log entry time 16:00:31 on December 5,2008

    Entry number 252191

    keyword=shift summary

    pi-, transverse
    08:00 in progress (30509,4211)
    08:24 end run 4211, q = 0.0395C
    08:25 start run 4212
    08:39 end run 4212, q = 0.0059C
    08:39 beam off ask for controlled access to check trigger supervisor problem and to change the LHRS polarity
    total charge 4171 -> 4212, q = 0.766C

    11:00 went into hall to find only Kaylan working,
    exit hall
    move target to empty
    ask for 2uA tuned beam
    ramped up Q2 to 1600A, Q3 to 1400A, set momentum to 2.35

    12:18 beam returns
    BeO check at 2 uA, positions at correct value (BPMA and BPMB).
    Did spot++ check. Put in He target, did raster scan.
    Target manual flip. Enable auto flip.

    begin production of pi+ transverse
    *** RICH HV was off for runs 4223-4225
    12:46 start run 4223
    12:59 end run 4223, q = 0.0064C
    13:00 start run 4224
    14:04 end run 4224, q = 0.0383C
    14:05 begin run 4225
    14:36 end run 4225, q = 0.0192C
    *** RICH HV off for previous three pi+ runs

    14:36 RICH HV turned on
    14:37 start run 4226
    15:39 end run 4226, q = 0.0384C
    15:40 start run 4227